OpenStudy (anonymous):

What two consecutive numbers is *radical sign*59 between? Lol I don't know how to do a radical sign so I improvised.. HELP!

OpenStudy (anonymous):

1^2=1, ..., 6^2 = 36, 7^2=49, 8^2 = 64, 9^2=81 Which two numbers is 59 between? {36, 49, 64, 81}? 59 is between 49 and 64 (ie 7^2 and 8^2) So sqrt(59) = 59^(1/2) is between sqrt(49) and sqrt(64) it is between 7 and 8

OpenStudy (anonymous):

Thanks!